# MoFO: Momentum-Filtered Optimizer for Mitigating Forgetting in LLM Fine-Tuning

### v3

@@ B.3 Proof of Theorem 2
Appendix C Supplemental Figures and Explanations
C.1 Supplementary Figures for Figure 1
C.2 Supplemental Explanation of Example 1
@@
For the $L_{2}$ regularization method, we add a regularization term $\lambda\|\theta-\theta_{\mathrm{pretrain}}\|^{2}_{2}$ to the fine-tuning loss to encourage the model to stay closer to the pre-trained state. As shown in Figure LABEL:appfig:example_landscape(b), the $L_{2}$ regularization approach remains closer to the pre-trained model, thereby achieving a smaller pre-training loss. However, since the fine-tuning objective is modified, Figure LABEL:appfig:example_landscape(a) shows that $L_{2}$ regularization does not reach the minimum of the fine-tuning loss.
@@
We note that HFT operates in a manner similar to Random BCD, which randomly selects a subset of coordinates (e.g., $\theta_{1}$ or $\theta_{2}$) to update. Figure LABEL:appfig:example_landscape further illustrates that both HFT and LoRA do not converge to minima as close to the pre-trained model as MoFO does, indicating that they may undergo higher levels of forgetting compared to MoFO.
@@
For LoRA, we make the following modelling in the landscape visualization example. The core principle of LoRA (Low-Rank Adaptation) involves approximating the original training space by a low-rank subspace. Since we consider a two-dimensional training space for visualizing the landscape, we set the rank of LoRA space to 1. Specifically, the parameters $\theta_{1}$ and $\theta_{2}$ exhibit a linear relationship. Given that the pretrained model is $(0,0)$, the parameters under LoRA are set to satisfy $\theta_{2}=\beta\theta_{1}$, where $\beta$ is a hyperparameter we set to 0.5. Figure LABEL:appfig:example_landscape shows that LoRA converges to a closer local minimum than Default FT.
Appendix D Implementation Details
D.1 Datasets for Fine-Tuning.
@@ Appendix D Implementation Details
MetaMathQA (Yu et al., 2024a). This dataset comprises 395K math question-answer pairs. Numerous studies indicate that LLMs significantly enhance performance metrics on mathematical benchmarks such as GSM8K after fine-tuning on this dataset. We randomly select 10% of this dataset for training LLMs, which includes 39.5K question-answer pairs.
@@
D.2 Evaluation Metrics for Instruction Fine-Tuning
@@
  • Common sense reasoning (CR): To measure the commonsense reasoning capabilities of LLMs, we employ the widely recognized benchmarks ARC-Challenge, ARC-Easy (Clark et al., 2018), and HellaSwag (Zellers et al., 2019), collectively referred to as the Commonsense benchmark. We use the average of their metrics as the evaluation, conducting assessments using the LM Eval Harness framework (Gao et al., 2023) and reporting the 0-shot accuracy based on the "acc_norm, none" metric.
@@
D.3 Hyperparameter Configurations
Instruction fine-tuning. In our instruction fine-tuning experiments, we follow the implementation of (Ivison et al., 2023). For instruction fine-tuning, we set the maximum sequence length to 1024, the global batch size to 128, and we train the model for 2 epochs. For the Llama-2-7B model, we use a learning rate of 2e-5 and 0 warm-up ratio, with a cosine decay learning rate scheduler. The learning rate is set to 2e-5 for fine-tuning both the Llama-2-7B-Chat model on the MetaMathQA dataset and the Gemma-2B-IT model, while a learning rate of 1e-5 is used for fine-tuning the Llama-2-7B-Chat model on the PMC-LLaMA-Instruct dataset; all these settings employ a warm-up ratio of 0.03 and a cosine decay learning rate scheduler. For LoRA, we set the learning rate as 1e-4. The other hyperparameters in the experiments are as follows.
@@ D.3 Hyperparameter Configurations
  • Update fraction of MoFO: $\alpha\%=15\%$.
@@
  • Update fraction of MoFO: $\alpha\%=15\%$.
@@
  • Update fraction of MoFO: $\alpha\%=15\%$.
@@
  • Update fraction of MoFO: $\alpha\%=5\%$.
@@
Hyperparameters in the Pareto comparison. To provide a comprehensive comparison, we explore various hyperparameter settings for $\lambda_{1}$, $\lambda_{2}$, LoRA’s rank, and the update fraction $\alpha\%$ in MoFO in Figure LABEL:pareto. Specifically, we set $\lambda_{1}$ as 1e-4, 1e-5, 1e-6, 1e-7, while $\lambda_{2}$ is set as 1e-2, 5e-3, 1e-3, 5e-4, and 1e-4. The update fraction $\alpha\%$ in MoFO is set as $5\%$, $10\%$, $15\%$, $20\%$, $40\%$, $80\%$. The rank of LoRA is set as 4, 16, 64, 256.
@@
D.4 More Explanation on the partitioning and Calculation of distance
Partitioning. We use the default partitioning scheme in PyTorch’s Transformer implementation. Different types of parameters within the Transformer, such as query (Q), key (K), value (V) weights for attention heads, and feed-forward network (FFN) weights, are divided into separate partitions. Notably, in the default PyTorch implementation, within a layer, the query (Q) weights of all attention heads are grouped into a single partition. The same applies to the key (K) and value (V) weights. Our momentum-based filtering mechanism is applied to each partition individually. A different parameter partition scheme, along with its corresponding experiments, is presented in Appendix LABEL:subapp:diff-partition.
@@ D.4 More Explanation on the partitioning and Calculation of distance
Appendix E Guideline for Setting $\alpha\%$
@@
    Grid Search: Perform a grid search over candidate values of $\alpha\%$ using this proxy subset.
@@
    Selection: Choose the $\alpha\%$ configuration that strikes a good balance between fine-tuning performance on the target dataset and preserving the model’s general capability.
To illustrate this procedure, we use Llama-2-7B as an example. We randomly sample 10% of the instances from the current training set of MetaMathQA dataset (39.5k) for fine-tuning, and then perform a grid search over $\alpha\%$ values of 5%, 10%, 15%, 20%, 40%, and 80%. As shown by the green line in Figure LABEL:fig:fraction_proxy(a), the fine-tuning performance is relatively stable across these values of $\alpha\%$. However, the green line in Figure LABEL:fig:fraction_proxy(b) indicates that $\alpha\%=15\%$ best preserves the model’s general capability. Therefore, we set $\alpha\%=15\%$.
Furthermore, by comparing the blue and green lines in Figure LABEL:fig:fraction_proxy, we observe that the trend of model performance with respect to $\alpha\%$ on the small proxy subset is consistent with the trend observed when fine-tuning on the full dataset. This implies that a small, randomly sampled subset is sufficient to guide the selection of a suitable $\alpha\%$.
Empirically, the optimal $\alpha\%$ often lies between 5% and 20%. A more fine-grained grid search within this range can be performed if needed. Designing more refined and efficient strategies for tuning $\alpha\%$ is left for future work.
Appendix F Additional Experiments on Instruction Fine-tuning
This section begins with a comparison of MoFO and baseline methods across additional datasets and models in F.1. In F.2, we explore the combination of LoRA and MoFO to assess their performance. Finally, in F.3, we compare MoFO with several algorithms designed to mitigate forgetting.
F.1 More Experimental Results in Instruction Fine-Tuning
@@
F.2 Experiment on the Combination of MoFO and LoRA

### v4

@@ A.3 Proof of Theorem 2 (Illustrative example: forgetting mitigation of MoFO)
A.4 Challenges and Potential Extensions of Theorem 1 to Nonsmooth Objectives
This subsection outlines challenges and possible future directions for extending our convergence analysis to nonsmooth objectives; a complete extension is left for future work.
Why the extension is challenging.
Our Theorem 1 relies on the standard $L$-smoothness assumption—i.e., the gradient of $\mathcal{L}$ is $L$-Lipschitz—to invoke the descent lemma and to derive an upper bound on $\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|$. When $\mathcal{L}$ is nonsmooth,
  1. (i)
    $\nabla\mathcal{L}(\theta)$ may not exist at nondifferentiable points, which requires working with generalized gradients (e.g., subgradients) rather than classical gradients in the derivation.
  2. (ii)
    The upper bound on $\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|$ in the proof of Theorem 1 (Appendix A.2) scales with the smoothness constant $L$; when $L$ is unbounded (or effectively very large), these inequalities become non-informative.
Possible extensions.
We outline below several plausible directions; full development is left for future work.
  1. (i)
    Subgradient analysis. Replace classical gradients with subgradients and assess convergence via a subgradient-based stationarity criterion.
  2. (ii)
    Smoothing. Introduce a family of smoothed surrogates $\mathcal{L}_{\mu}$}, analyze the MoFO algorithm under these surrogates to obtain $\mu$-dependent bounds, and then let $\mu\downarrow 0$ to recover convergence results for the original loss $\mathcal{L}$.
  3. (iii)
    Algorithmic modifications. Incorporate techniques such as gradient clipping or stochastic subgradient steps (Xiao et al., 2024), and then analyze the convergence of the modified algorithm.
Appendix B Supplemental Figures and Explanations
B.1 Reason for using RedPajama to approximate LLaMA-2’s training data
We note that original LLaMA-2 training dataset has not been publicly released. Thus, we can only rely on public datasets to approximate LLaMA-2’s original training data.
RedPajama project was explicitly designed as an open-source reproduction of the LLaMA training dataset (Weber et al., 2024; Together, 2023). It closely mirrors the data sources outlined in the original LLaMA paper and adopts similar strategies for data collection, mixture, and preprocessing. We believe it serves as a reasonable proxy for approximating LLaMA-2’s training dataset.
B.2 Supplementary Figures for Figure 2(a)
[…]
B.3 Supplementary Experiments on the Correlation between Distance and Forgetting
(b) Average accuracy on the MMLU benchmark (measuring preservation of factual knowledge) for Llama-2-7B after fine-tuning on MetaMathQA with Adam, Lion, and MoFO. Building on Figure 2(b)(b), we add points for runs where Llama-2-7B was trained for 0.1, 0.2, and >3 epochs using both Adam and MoFO. The marker size encodes the number of training epochs (larger means more epochs). Red arrows indicate the points obtained after exactly 1 epoch for each optimizer.
In this subsection, we augment Figure 2(b)(b) by probing the relationship between a model’s parameter distance from its pre-trained state and evaluation accuracy under additional training budgets and optimizers. Concretely, under the same settings as Figure 2(b)(b), we add runs at 0.1 and 0.2 epochs and extend training beyond 3 epochs, using Adam and MoFO. We report results on MMLU (as in the main text, measuring preservation of factual knowledge) and newly include HumanEval (measuring preservation of code-generation ability). Since the fine-tuning task is math, both can serve as forgetting mitigation metrics. The corresponding scatter plots are shown in Figure 8(b) (MMLU) and Figure 8(c) (HumanEval). When examining the points for each optimizer separately, we make the following observations:
  • Observation 1 (sufficient training). Once training exceeds approximately 1 epoch, MMLU and HumanEval scores show a consistent strong negative correlation with the parameter distance to the pre-trained state.
  • Observation 2 (early training). For Adam or MoFO at less than 1 epoch, we may observe a mild correlation with the parameter distance to the pre-trained state. The correlation may be unstable and can be positive or negative.
We speculate that this short-lived positive trend may be related to benchmark alignment. The MMLU benchmark (used to measure the preservation of factual knowledge) might share partial overlap with the patterns of our math fine-tuning task (measured by GSM8K benchmark); in the early training steps, the model might incidentally acquire features that also benefit MMLU, resulting in a temporary gain. By contrast, as for HumanEval in Figure 8(c), which measures code generation and differs from our math fine-tuning in both domain and output format, it may exhibit an unstable correlation whose sign can be either negative or positive. Another possible factor could stochasticity, since at less than 1 epoch the dataset has not yet been fully traversed.
Overall, the negative correlation becomes clear and consistent after sufficient training; while the early training presents a mild, benchmark-dependent relationship.
(c) Scores on HumanEval benchmark for Llama-2-7B after fine-tuning on MetaMathQA with Adam, Lion, and MoFO. Building on Figure 2(b)(b), we add points for runs where Llama-2-7B was trained for 0.1, 0.2, and >3 epochs using both Adam and MoFO. The marker size encodes the number of training epochs (larger means more epochs). Red arrows indicate the points obtained after exactly 1 epoch for each optimizer.
In addition, we emphasize an empirical point: the negative relationship between parameter distance and the preservation of pre-trained knowledge is evident across optimizers. In Figure 2(b)(b), 8(b), and 8(c), the parameter distances roughly follow the ordering Lion > Adam > MoFO, whereas the scores measuring preservation of pre-trained knowledge follow the inverse ordering MoFO > Adam > Lion. For a broader comparison, we evaluate five optimizers—MoFO, NAdam (Dozat, 2016), Adam, RMSProp (Tieleman, 2012), and Lion. After two epochs of training, we report (i) their parameter distance to the pre-trained state and (ii) their forgetting-mitigation performance on MMLU and HumanEval, shown in Figure 8(d)(a) and Figure 8(d)(b), respectively. The results show a consistent negative correlation between distance and performance. Notably, MoFO remains closer to the pre-trained state and achieves higher scores compared with the other optimizers. Therefore, this cross-optimizer rank-order correlation is sufficient to motivate our algorithm design: favor optimizers that converge closer to the pre-trained state, so as to better preserve the pre-trained knowledge.
(d) (a) MMLU accuracies and (b) HumanEval scores of Llama-2-7B after fine-tuning on the MetaMathQA dataset using different optimizers. All experiments are run for 2 epochs.
B.4 Supplemental Explanation of Example 1
(e) The loss landscapes of the example. We plot the landscapes on (a) the fine-tuning loss and (b) the pre-training loss. The color bar indicates the loss value—(a) fine-tuning loss and (b) pre-training loss. A logarithmic scale is applied to the loss values for better visualization. In Example 1, MoFO converges to a minimum closest to the pre-trained model, with a low pre-training loss.
@@ B.4 Supplemental Explanation of Example 1
For the $L_{2}$ regularization method, we add a regularization term $\lambda\|\theta-\theta_{\mathrm{pretrain}}\|^{2}_{2}$ to the fine-tuning loss to encourage the model to stay closer to the pre-trained state. As shown in Figure 8(e)(b), the $L_{2}$ regularization approach remains closer to the pre-trained model, thereby achieving a smaller pre-training loss. However, since the fine-tuning objective is modified, Figure 8(e)(a) shows that $L_{2}$ regularization does not reach the minimum of the fine-tuning loss.
@@
We note that HFT operates in a manner similar to Random BCD, which randomly selects a subset of coordinates (e.g., $\theta_{1}$ or $\theta_{2}$) to update. Figure 8(e) further illustrates that both HFT and LoRA do not converge to minima as close to the pre-trained model as MoFO does, indicating that they may undergo higher levels of forgetting compared to MoFO.
@@
For LoRA, we make the following modelling in the landscape visualization example. The core principle of LoRA (Low-Rank Adaptation) involves approximating the original training space by a low-rank subspace. Since we consider a two-dimensional training space for visualizing the landscape, we set the rank of LoRA space to 1. Specifically, the parameters $\theta_{1}$ and $\theta_{2}$ exhibit a linear relationship. Given that the pre-trained model is $(0,0)$, the parameters under LoRA are set to satisfy $\theta_{2}=\beta\theta_{1}$, where $\beta$ is a hyperparameter we set to 0.5. Figure 8(e) shows that LoRA converges to a closer local minimum than Default FT.
B.5 Synthetic Experiment for Example 1
In this subsection, we conduct a synthetic experiment to provide a more concrete illustration of Example 1. Specifically, we set the parameter dimension to $d=10$, with parameters $\theta=(\theta_{1},\dots,\theta_{10})$. The pretraining loss is defined as the squared $L_{2}$ norm of the parameters: $L_{\rm{pretrain}}(\theta)=\frac{1}{2}\|\theta\|_{2}^{2}$, with the pre-trained model given by $\theta_{\rm pretrain}=(0,0,\dots,0)$. Starting from the pre-trained model, we optimize the parameters with respect to the fine-tuning loss $\mathcal{L}(\theta)=\prod_{i=1}^{d}(a_{i}\theta_{i}-b_{i})^{2}$, where $a_{i},b_{i}>0$ for any $1\leq i\leq d$. The coefficients $a_{i}$ and $b_{i}$ are sampled from a standard normal distribution; to ensure positivity, we take their absolute values and add 0.3 and 0.1, respectively.
In this experiment, we compare two optimizers: Adam and MoFO. For each, we perform a grid search for the optimal learning rate over the set $\{10^{-2},10^{-3},10^{-4}\}$. We consider the fine-tuning process to have converged to a minimum when the fine-tuning loss drops below $10^{-8}$ within 10000 iterations. At convergence, we record two metrics: the Euclidean distance from the fine-tuned model to the original pre-trained model, and the value of the pretraining loss. The entire experiment is repeated across three different random seeds for robustness.
As presented in Table 4, the results indicate that the minimum found by MoFO is at roughly half the distance from the pre-trained model compared to the one found by Adam. Concurrently, MoFO achieves a lower pre-training loss. These findings provide strong evidence that MoFO can converge to minima closer to the pre-trained model in Example 1, thereby supporting Theorem 2.
Table 4: The Euclidean distance from the fine-tuned model to the original pre-trained model, and the pretraining loss of parameters after optimizing the fine-tuning loss using Adam and MoFO. The results show that MoFO finds a fine-tuning minimum that is closer to the pre-trained model compared to the one found by Adam.
Adam MoFO
Distance to pre-trained model 0.549 0.275
Pre-training loss 0.151 0.038
Appendix C Implementation Details
C.1 Datasets for Fine-Tuning.
@@ Appendix C Implementation Details
MetaMathQA (Yu et al., 2024b). This dataset comprises 395K math question-answer pairs. Numerous studies indicate that LLMs significantly enhance performance metrics on mathematical benchmarks such as GSM8K after fine-tuning on this dataset. We randomly select 10% of this dataset for training LLMs, which includes 39.5K question-answer pairs.
@@
Magicoder-Evol-Instruct (Wei et al., 2023). This dataset comprises 110K instruction-response pairs related to coding, and it is decontaminated and redistributed from the Evol-CodeAlpaca-V1 dataset (Luo et al., 2023b) We randomly select 39.5K question-answer pairs from this dataset to fine-tune LLM and enhance their coding capability.
@@
C.2 Evaluation Metrics for Instruction Fine-Tuning
@@
  • Common sense reasoning (CR): To measure the commonsense reasoning capabilities of LLMs, we employ the widely recognized benchmarks ARC-Challenge (ARC-C), ARC-Easy (ARC-E) (Clark et al., 2018), and HellaSwag (Zellers et al., 2019), collectively referred to as the Commonsense benchmark. We use the average of their metrics as the evaluation, conducting assessments using the LM Eval Harness framework (Gao et al., 2023) and reporting the 0-shot accuracy based on the "acc_norm, none" metric.
@@
C.3 Hyperparameter Configurations
Instruction fine-tuning. In our instruction fine-tuning experiments, we follow the implementation of Ivison et al. (2023). For instruction fine-tuning, we set the maximum sequence length to 1024, the global batch size to 128, and we train the model for 2 epochs. For the Llama-2-7B model, we use a learning rate of 2e-5 and 0 warm-up ratio, with a cosine decay learning rate scheduler. The learning rate is set to 2e-5 for fine-tuning both the Llama-2-7B-Chat model on the MetaMathQA dataset and the Gemma-2B-IT model, while a learning rate of 1e-5 is used for fine-tuning the Llama-2-7B-Chat model on the PMC-LLaMA-Instruct dataset; all these settings employ a warm-up ratio of 0.03 and a cosine decay learning rate scheduler. For LoRA, we set the learning rate as 1e-4. The other hyperparameters in the experiments are as follows.
@@ C.3 Hyperparameter Configurations
  • Update fraction of MoFO: $\alpha=15\%$.
@@
  • Update fraction of MoFO: $\alpha=15\%$.
@@
  • Update fraction of MoFO: $\alpha=15\%$.
@@
  • Update fraction of MoFO: $\alpha=5\%$.
@@
Fine-tuning Llama-2-7B-Chat on Magicoder-Evol-Instruct.
  • Learning rate: 1e-5.
  • Update fraction of MoFO: $\alpha=20\%$.
Hyperparameters in the Pareto comparison. To provide a comprehensive comparison, we explore various hyperparameter settings for $\lambda_{1}$, $\lambda_{2}$, LoRA’s rank, and the update fraction $\alpha$ in MoFO in Figure 5(a). Specifically, we set $\lambda_{1}$ as 1e-4, 1e-5, 1e-6, 1e-7, while $\lambda_{2}$ is set as 1e-2, 5e-3, 1e-3, 5e-4, and 1e-4. The update fraction $\alpha$ in MoFO is set as $5\%$, $10\%$, $15\%$, $20\%$, $40\%$, $80\%$. The rank of LoRA is set as 4, 16, 64, 256.
@@
C.4 More Explanation on the partitioning and Calculation of distance
Partitioning. We use the default partitioning scheme in PyTorch’s Transformer implementation. Different types of parameters within the Transformer, such as query (Q), key (K), value (V) weights for attention heads, and feed-forward network (FFN) weights, are divided into separate partitions. Notably, in the default PyTorch implementation, within a layer, the query (Q) weights of all attention heads are grouped into a single partition. The same applies to the key (K) and value (V) weights. Our momentum-based filtering mechanism is applied to each partition individually. A different parameter partition scheme, along with its corresponding experiments, is presented in Appendix F.6.
@@ C.4 More Explanation on the partitioning and Calculation of distance
Appendix D Guideline for Setting $\alpha$
@@
    Grid Search: Perform a grid search over candidate values of $\alpha$ using this proxy subset.
@@
    Selection: Choose the $\alpha$ configuration that strikes a good balance between fine-tuning performance on the target dataset and preserving the model’s general capability.
To illustrate this procedure, we use Llama-2-7B as an example. We randomly sample 10% of the instances from the current training set of MetaMathQA dataset (39.5k) for fine-tuning, and then perform a grid search over $\alpha$ values of 5%, 10%, 15%, 20%, 40%, and 80%. As shown by the green line in Figure 10(a)(a), the fine-tuning performance is relatively stable across these values of $\alpha$. However, the green line in Figure 10(a)(b) indicates that $\alpha=15\%$ best preserves the model’s general capability. Therefore, we set $\alpha=15\%$.
(a) (a) Fine-tuning performance: Accuracy on the GSM8K math reasoning task for LLMs of different sizes, fine-tuned via MoFO with varying update fractions ($\alpha$). (b) Preservation of pre-training knowledge: Average accuracy changes on MMLU, HumanEval, and commonsense reasoning benchmarks relative to the original pre-trained LLMs, illustrating how much pre-training knowledge is retained. All results are obtained by fine-tuning Llama2-7B on MetaMathQA and its proxy subset. The performance trends under different update fractions on the proxy subset align with those observed on the full dataset.
Furthermore, by comparing the blue and green lines in Figure 10(a), we observe that the trend of model performance with respect to $\alpha$ on the small proxy subset is consistent with the trend observed when fine-tuning on the full dataset. This implies that a small, randomly sampled subset is sufficient to guide the selection of a suitable $\alpha$.
Empirically, the optimal $\alpha$ often lies between 5% and 20%. A more fine-grained grid search within this range can be performed if needed. Designing more refined and efficient strategies for tuning $\alpha$ is left for future work.
Appendix E Additional Experiments on Instruction Fine-tuning
This section begins with a comparison of MoFO and baseline methods across additional datasets and models in E.1. In E.2, we explore the combination of LoRA and MoFO to assess their performance. Finally, in E.3, we compare MoFO with several algorithms designed to mitigate forgetting.
E.1 More Experimental Results in Instruction Fine-Tuning
@@
Method MedQ General Capability
CR IFEval HumanEval Avg.
Llama-2-7B-Chat 49.8 65.6 41.4 24.3 43.8
Default FT 54.3 64.6 32.1 20.6 39.1
HFT 54.4 65.2 33.5 23.1 40.6
LoRA 54.2 64.4 33.9 23.5 40.6
MoFO 54.3 65.6 38.6 25.0 43.1
@@
Method GSM8K General Capability
CR IFeval HumanEval Avg.
Gemma-2B-IT 11.4 57.6 33.6 31.5 40.9
Default FT 42.0 52.1 24.3 20.6 32.3
HFT 41.5 53.9 24.1 21.2 33.1
LoRA 40.6 54.4 26.1 29.8 36.8
MoFO 42.1 55.0 28.7 29.1 37.6
@@
Method GSM8K General Capability
CR IFeval HumanEval Avg.
Llama-2-7B-Chat 13.7 65.6 41.4 24.3 43.8
Default FT 48.4 62.8 30.7 15.6 36.4
HFT 46.9 63.4 31.8 20.0 38.4
LoRA 45.3 63.9 35.6 21.0 40.2
MoFO 47.1 64.0 37.1 21.7 40.9
@@
Table 8: The performance of the fine-tuning task (coding), measured by HumanEval, and the general capability scores of Llama-2-7B-Chat after fine-tuning on the Magicoder-Evol-Instruct dataset. Here we choose the three benchmarks exhibiting the most significant forgetting. We set the rank of LoRA as 256, and $\alpha$ of MoFO is set as 20%. The results show that MoFO achieves comparable performance in the fine-tuning task, while mitigating forgetting of general capabilities. Bold values denote the best results among these methods.
Method HumanEval General Capability
ARC-E ARC-C IFEval Avg.
Llama-2-7B-Chat 24.2 74.5 46.3 41.1 54.0
Default FT 56.2 71.2 45.2 33.5 50.0
HFT 50.7 71.5 45.6 36.3 51.1
LoRA 48.6 72.1 45.1 33.4 50.2
MoFO 53.3 72.1 46.0 36.1 51.4
Results of fine-tuning Llama-2-7B-Chat on Magicoder-Evol-Instruct.
We also fine-tune the Llama-2-7B-Chat on the Magicoder-Evol-Instruct dataset. We use ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) scores to measure general capability. The results in Table 8 demonstrate that our approach outperforms the baselines in the fine-tuning tasks and exhibits less forgetting compared to baseline methods.
@@
E.2 Experiment on the Combination of MoFO and LoRA